OpenStudy (anonymous):

log2^n=1/4log2^16+1/2log2^49?

OpenStudy (ivancsc1996):

You can drop the n since it is a logarith and then solve:\[\log(2^{n})=nlog2 \rightarrow n=\frac{ \frac{ 1 }{ 4 }\log(2^{16})+\frac{ 1 }{ 2 }\log (2^{49}) }{ \log2 }=\frac{ 4+\frac{ 49 }{ 2 } }{ 1 }=\]

OpenStudy (anonymous):

Thanx!